What is dielectric constant of a medium.p plz answer to me as soon as possible.

Grade:12

1 Answers

Susmita
5 years ago
Dielectric constant is the electrical permittivity of a substance divided by permittivity of free space.
K=\frac{\epsilon}{\epsilon_{0}}=\frac{D}{\epsilon_{0} E}=\frac{\epsilon_{0} E+P}{\epsilon_{0} E}=1+\frac{P}{\epsilon_{0}E}
So basically it is a measure of how much a material deviates from conducting property.Because polarization P occur in bound charges or insulating materials.
It means when we place a insulating material under an electric field E,it allows the electric field to polarize the charges inside it or in other words,allows the electric field to penetrate it.
Whereas,conductors do not allow electric field to penetrate it,proof is that charges reside outside the surface.So there is no electric field inside a conductor.So k=1+P/0=infinite.
So K is also a measure of how much resistance a substance give to an electric field to penetrate it.
As the conductivity of insulator increases on increasing temperature,and that of a conductor decreses,so K also changes with temperature.